# Is this the real life? Is this just fantasy? The Misleading Success of Simulating Social Interactions With LLMs

### v3

@@ 3.1 The Unified Framework for Simulation
We build on the Sotopia framework (Zhou et al., 2024), in which 40 unique characters with relationships interact in 90 diverse social scenarios. We then simulate social interactions across various setting under a unified framework. Sotopia is built on Agents mode interactions natively, allowing agents to generate utterances (e.g., Ben said: “how are you?”), non-verbal communication (e.g., Ben smiled), and actions (e.g., Ben moved to the room). We add these other modes of simulations for comparison.
@@ 4.3 RQ3: Training on Script simulations results in selective improvements
We find that learning from Script simulations enhances naturalness but not necessarily for goal completion in cooperative scenarios. As shown in Figure 5, the finetuned Agents model achieves a higher goal completion rate than the original GPT-3.5 in the Agents mode. However, the finetuned model still performs significantly worse than the Script mode. In cooperative scenarios (§3.3), the finetuned model barely improves, where seeking common information is critical to the task’s success. As shown in Figure 6, the finetuned model struggles to complete the social goals in the Agents mode by following the strategies of Script simulations. In contrast, the finetuned model shows a relatively large improvement in the competitive scenarios. Also, finetuning significantly improves Agents’s naturalness, as evidenced by the finetuned model’s naturalness is not different from the Script mode according to human evaluation. [^10] Please see Appendix E for more details.
@@ 4.4 RQ4: Script simulations can be biased
To illustrate the limitations of Script mode, we explore task-specific metrics to understand why finetuning improves for competitive but not cooperative scenarios. For the competitive scenarios, we measure how often the interaction ends in an agreement as a proxy for the agreeableness of the interaction style. Specifically, we calculate the percentage of the interactions that end in a successful purchase in the Craigslist task. [^11] We use GPT-4 to determine whether the interaction ends in an agreement. Please refer to the Appendix H for the details. We find that the Script simulations reach a deal in 94% of the interactions, while Agents simulations only reach a deal in 30% of the interactions. Finetuning the model increases the percentage to 93%, which indicates that models can easily follow this overly agreeable style from Script simulations.
@@ Acknowledgements
First of all, we thank our graduate student annotators for helping us with judging the naturalness of the simulations. We thank Hao Zhu and Daniel Fried for their discussions and feedback. We also thank OpenAI and Together AI for providing credits for running the models in this work. TE acknowledges support from the GEM consortium and the National Science Foundation Graduate Research Fellowship under Grant No. 1745302.
@@ Appendix C Example Code Snippets for Determining Simulation Modes
We provide example code snippets for determining the simulation modes in Park et al. (2023). The code is from the official Github repo of Park et al. (2023).  [^14] https://github.com/joonspk-research/generative_agents/blob/fe05a71d3e4ed7d10bf68aa4eda6dd995ec070f4/reverie/backend_server/persona/cognitive_modules/plan.py#L277

### v4

@@ 3.1 The Unified Framework for Simulation
We build on the Sotopia framework (Zhou et al., 2024), in which 40 unique characters with relationships interact in 90 diverse social scenarios. A social task in Sotopia involves a scenario, two character profiles, and their respective private social goals for the interaction. During an episode, the two agents, whether AI or human, role-play the characters to accomplish their social goals. Agents are allowed to generate utterances (e.g., Ben said: “how are you?”), non-verbal communication (e.g., Ben smiled), and actions (e.g., Ben moved to the room).
Sotopia primarily focuses on general social interactions between agents, where each agent has distinct social goals and different information about the other (Agents). To provide a broader comparison, we introduce additional simulation modes. These various settings are then simulated under a unified framework to analyze the social interactions comprehensively.
@@ 4.3 RQ3: Training on Script simulations results in selective improvements
We find that learning from Script simulations enhances naturalness but not necessarily for goal completion in cooperative scenarios. As shown in Figure 5, the finetuned Agents model achieves a higher goal completion rate than the original GPT-3.5 in the Agents mode. However, the finetuned model still performs significantly worse than the Script mode. In cooperative scenarios (§3.3), the finetuned model barely improves, where seeking common information is critical to the task’s success. As shown in Figure 6, the finetuned model struggles to complete the social goals in the Agents mode by following the strategies of Script simulations. In the Script simulations, the model has omniscient knowledge of each agents’ goals and information (e.g., the mutual friend’s name in the MutualFriends scenario), therefore, it can easily complete the social goals by exploiting the information (i.e., the agent spits out the mutual friend’s name accurately). However, such strategies are not applicable in the Agents mode, where the model does not have access to the other agents’ goals and information.
In contrast, the finetuned model shows a relatively large improvement in the competitive scenarios. However, this does not necessarily mean that the finetuned model is improving its negotiation skills through learning the demonstrations from the Script simulations. As in the competitive scenarios, the agents can be overly agreeable to reach an agreement without actually negotiating with each other. Meanwhile, finetuning significantly improves Agents’s naturalness, as evidenced by the finetuned model’s naturalness is not different from the Script mode according to human evaluation. This suggests that the finetuned model learns the interaction style from the Script simulations.  [^10] Please see Appendix E for more details.
@@ 4.4 RQ4: Script simulations can be biased
To illustrate the limitations of Script mode, we explore task-specific metrics to understand why finetuning improves for competitive but not cooperative scenarios. For the competitive scenarios, we measure how often the interaction ends in an agreement as a proxy for the agreeableness of the interaction style. Specifically, we calculate the percentage of the interactions that end in a successful purchase in the Craigslist task. [^11] We use GPT-4 to determine whether the interaction ends in an agreement. Please refer to the Appendix H for the details. We find that the Script simulations reach a deal in 94% of the interactions, while Agents simulations only reach a deal in 30% of the interactions. Finetuning the model increases the percentage to 93%, which indicates that models can easily follow this overly agreeable style from Script simulations. This explains the large improvement of finetuning on Script simulations for competitive scenarios, which is not due to learning the negotiation skills but more likely due to learning the interaction style from the Script simulations.
@@ Acknowledgements
First of all, we thank our graduate student annotators for helping us with judging the naturalness of the simulations. We thank Hao Zhu, Daniel Fried, Carolyn Rosé, Kaitlyn Zhou and Jenny Liang for their discussions and feedback. We also thank OpenAI and Together AI for providing credits for running the models in this work. TE acknowledges support from the GEM consortium and the National Science Foundation Graduate Research Fellowship under Grant No. 1745302. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410.
@@ Appendix C Example Code Snippets for Determining Simulation Modes
We provide example code snippets for determining the simulation modes in Park et al. (2023). The code is from the official Github repo of Park et al. (2023).